OpenStudy (anonymous):

how do I find the equation of the tangent line at y=(2x+2)e^x given the point p=2,6e^2

OpenStudy (tkhunny):

What do you get when you evaluate the 1st derivative for a given value of x?

OpenStudy (anonymous):

the slope, and i got 2e^x(x+2) but i dont know what to do next

OpenStudy (tkhunny):

The slope of what?

OpenStudy (anonymous):

the equation of the tan line?

OpenStudy (tkhunny):

Very good. If the derivative is correct, you are ready to proceed. You have a point. You have a slope. You have the Point-Slope form of a line: \((y-y_{0}) = m(x-x_{0})\) Go!

OpenStudy (anonymous):

my answer has to be in y=mx+b form, and I cant seem to simplify it :( I have 4 options to choose from and my answer isnt matching up

OpenStudy (tkhunny):

x = 2 \(y = (2(2)+2)\cdot e^{2} = 6\cdot e^{2}\) You have a point: \((2,6\cdot e^{2})\) As you correctly calculated: \(y' = 2(x+2)\cdot e^{x}\) Again, with x = 2 \(y'(2) = 2(2+2)\cdot e^{2} = 8\cdot e^{2}\) You have a slope: \(8\cdot e^{2}\) Now, we apply the Point-Slope form of the line!

OpenStudy (tkhunny):

Great! Let's see what you get and how you get it.

OpenStudy (anonymous):

ohhhhhhhhh you plug the 2 back into the slope !! i was trying to figure out an answer with 2e^x(x+2) and it just wasn't working, so does 8e^2(x) - 10e^2 make sense?

OpenStudy (tkhunny):

\((y - 6e^{2}) = 8e^{2}(x-2)\) \(y - 6e^{2} = 8e^{2}(x) - 16e^{2}\) \(y = 8e^{2}(x) - 10e^{2}\) It's not pretty, but there it is! Good work.

OpenStudy (tkhunny):

That's really awkward-looking. I think I would prefer \(8xe^{2}\) rather than \(8e^{2}x\).

OpenStudy (anonymous):

no joke 8e^2x was how it was written in my text book as a possible answer, thanks so much though